OpenStudy (hellomynameiskam):

Solve. - s / 3 > 6 A. s ≥ –2 B. s ≤ –2 C. s ≥ –18 D. s ≤ –18

OpenStudy (kendricklamar2014):

Is the question: \[\frac{ -s }{ 3 } \ge 6\] or \[\frac{ -s }{ 3 } > 6\]